# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v4

@@ D.1.1 Alaa Synthetic Data Experiments
The first synthetic data experiments follows the setup proposed by Alaa et al. [15]. Two setups are used: A and B. Each setup uses 5000 samples and 10 covariates. The experiments are performed using 100 simulations, each using a different random seed and using the $\alpha$ value of $0.10$. Additionally, for our proposed model, we evaluate the probabilistic calibration of the predictive distribution. The CCT and CMC models use normalized split WCPS. The results are shown in Figure  2(a) and 2(b).
@@
This data generation process entails both Setup A and Setup B. For Setup A, where $\gamma=1$, represents an experiment where the treatment has no effect, while for Setup B, where $\gamma=0$, represents a heterogeneous treatment effect. X is sampled from a uniform distribution, $d$ is the number of covariates, $\tau$ represents the treatment effect, $\gamma$ is the treatment heterogeneity, $\pi$ is the propensity, $\epsilon_{0}$ and $\epsilon_{1}$ are Gaussian noise terms, and $T$ is the treatment assignment.
$X\sim uniform([0,1],size=n*d)$ (12)
@@
$\epsilon_{1}\sim\mathcal{N}(0,1)$ (16)
$\epsilon_{0}\sim\mathcal{N}(0,1)$ (17)
$y_{0}=\tau_{0}+\epsilon_{0}$ (18)
$y_{1}=\tau+\epsilon_{1}$ (19)
@@
$T\sim(uniform(0,1)<\pi)$ (20)
$y=T\cdot y_{1}+(1-T)\cdot y_{0}$ (21)
D.1.2 Nie and Wager Synthetic Data Experiments
The second synthetic data experiments use the setups of Nie and Wager [5]. Four setups are used: A, B, C, and D. Each setup uses 5000 samples and is split into a training and testing set using a 50% split. Every setup is simulated with uncorrelated noise terms, uses five covariates, and a standard deviation of the error term of 1. The same approach is taken here as in Section D.1.1 with 100 simulations, each using different random seeds and using the same $\alpha$ values. The CCT and CMC models also use normalized split WCPS. The results are shown in Figure 8,  9, and  10.
Besides the aforementioned experiments with the data-generating processes proposed by Nie and Wager [5], we also perform a noise dependency impact experiment with it. This experiment will investigate the impact of changing the dependency between the noise variables, $\epsilon_{0}$ and $\epsilon_{1}$, of $y_{0}$ and $y_{1}$, respectively, on the different conformal frameworks. The only difference with the previous experiment is that the noise correlation coefficient $c$ changes from $-1$ to $1$ in increments of $0.1$. The results are shown in Figure 3.
@@ D.1.2 Nie and Wager Synthetic Data Experiments
Setup A simulates a difficult nuisance component and an easy treatment effect. $p$ is the number of covariates and $n$ the number of samples. X is created as an independent variable, $n$ rows and $p$ columns thus consisting of $[x_{0},x_{1},...,x_{p}]$ each with length $n$. The other variables are the expected outcome $b$, the propensity of receiving treatment $\pi$, $\tau$ the ITE, $w$ the treatment assignment, $y_{0}$ outcome for control, $y_{1}$ for treatment, $\epsilon_{0}$ noise variable for control, $\epsilon_{1}$ noise variable for treatment, and $y$ the true outcome, all with length $n$.
$X\sim uniform([0,1],size=n*p)$ (22)
$b=sin(\pi\cdot x_{0}\cdot x_{1})+2(x_{2}-0.5)^{2}+x_{3}+0.5x_{4}$ (23)
$\pi=max([0.1,0.1,...,0.1],min(sin(\pi\cdot x_{0}\cdot x_{1}),[0.9,0.9,...,0.9])$ (24)
@@
$w\sim Binomial(1,e)$ (26)
$\epsilon_{0}\sim\mathcal{N}(0,\sigma^{2})$ (27)
$\epsilon_{1}=c\cdot\epsilon_{0}+(1-c)\cdot\mathcal{N}(0,\sigma^{2})$ (28)
$y_{0}=b-0.5\tau+\epsilon_{0}$ (29)
$y_{1}=b+0.5\tau+\epsilon_{1}$ (30)
$y=y_{0}(1-w)+y_{1}\cdot w$ (31)
@@
Setup D mimics a situation with unrelated treatment and control groups. All other variables are calculated similarly as in setup B except for $b,e,$ and $\tau$.
@@
D.2 Semi-Synthetic Experiment
The learners are also evaluated on semi-synthetic datasets following the experiments of Alaa et al. [15]. Three semi-synthetic datasets are used: IHDP [10], NLSM [33], and ACIC2016[31] dataset. The simulations run 100 times with $\alpha=0.1$, 100 MC samples, and with split WCPS using the residual as a conformity measure. Since these datasets do not provide propensity scores, we perform logistic regression to estimate it. The results are found in Figure 4 and  11.
D.2.1 IHDP
The Infant Health and Development Program (IHDP) is a randomized controlled trial to evaluate the effect of home visits from physicians on the cognitive test scores of premature infants [10]. This dataset has 25 covariates with 6 continuous and 19 binary. Using this dataset, a semi-synthetic benchmark was created analogous to the study by Hill, using setup B of their work [10]. The used datasets are provided by Shalit et al. [34]. The final dataset consists of 747 samples, 139 are treated samples and 608 of them are control samples. The data generation process of the semi-synthetic data is as follows:
$Y_{0}\sim\mathcal{N}(exp(X+W)\beta,1)$ (42)
$Y_{1}\sim\mathcal{N}(exp(X\beta)-\omega,1)$ (43)
@@ D.2.1 IHDP
respectively. X represents the covariates in the IHDP dataset and W is an offset matrix. The same treatment variable is kept as in the original IHDP study.
D.2.2 NLSM
The National Study of Learning Mindsets (NLSM) is a large-scale randomized trial to investigate the effect of behavioural intervention on the academic performance of students [33]. Using the original NLSM dataset, new data is synthetically generated that follow the same distributions as the NLSM dataset as elaborated in Carvalho et al. [33], Alaa et al. [15]. The resulting dataset contains 10000 data points and 11 covariates ($X=[x_{0},x_{1},...,x_{10}]$). The data generation process for the semi-synthetic data is as follows:
@@ D.2.2 NLSM
$\tau=0.228+0.05\cdot\mathcal{I}(x_{6}<0.07)-0.05\cdot\mathcal{I}(x_{7}<-0.69)-0.08\cdot\mathcal{I}(x_{4}\in[1,13,14]+\gamma_{x_{0}-1}$ (45)
with $\mathcal{I}$ the indicator function and $\gamma_{x_{0}-1}$ the $(x_{0}-1)$th element of an array of Gaussian samples with size 76. Then the outcome is defined as follows:
$Y=W(Y-\tau)+(1-W)*\tau$ (46)
@@
D.2.3 ACIC2016
The dataset from the 2016 Atlantic Causal Inference Conference was created for a causal inference competition [31]. It includes 4,802 observations with 58 covariates (3 categorical, 5 binary, 27 counts, and 23 continuous). The competition used these covariates to generate 77 different settings, each with varying outcome surface complexity, degrees of confounding, overlap, and treatment effect heterogeneity. We standardized all covariates and, for each setting, conducted 10 simulation runs, with a test set comprising 20% of the data, randomly generated, following the methodology of Curth and van der Schaar [35], Alaa et al. [15].
Appendix E Experiments Results
Figure 9: Simulation results for setups A, B, C, and D from Nie and Wager [5] evaluating probabilistic calibration (PIT plot) of CCT- and CMC T-learner with WCPS.
Figure 10: Simulation results for setups A, B, C, and D from Nie and Wager [5] evaluating probabilistic calibration (PIT plot) of CCT- and CMC T-learner with unweighted CPS.

### v5

@@ G.2 Synthetic experiments
The first synthetic data experiments follows the setup proposed by Alaa et al. [2023]. Two setups are used: A and B. Each setup uses 5000 samples and 10 covariates. The experiments are performed using 100 simulations, each using a different random seed and using the $\alpha$ value of $0.10$. Additionally, for our proposed model, we evaluate the probabilistic calibration of the predictive distribution. The CCT and CMC models use normalized split WCPS. The results are shown in Figure 3, 9, and 10.
@@
This data generation process entails both Setup A and Setup B. For Setup A, where $\gamma=1$, represents an experiment where the treatment has no effect, while for Setup B, where $\gamma=0$, represents a heterogeneous treatment effect. X is sampled from a uniform distribution, $d$ is the number of covariates, $\tau$ represents the treatment effect, $\gamma$ is the treatment heterogeneity, $\pi$ is the propensity, $\epsilon(0)$ and $\epsilon(1)$ are Gaussian noise terms, and $W$ is the treatment assignment.
$X\sim Uniform([0,1],size=(n,d))$ (26)
@@
$\epsilon(1)\sim\mathcal{N}(0,1)$ (30)
$\epsilon(0)\sim\mathcal{N}(0,1)$ (31)
$Y(0)=\tau_{0}+\epsilon(0)$ (32)
$Y(1)=\tau+\epsilon(1)$ (33)
@@
$W\sim(Uniform(0,1)<\pi)$ (34)
$y=WY(1)+(1-W)Y(0)$ (35)
G.2.2 Nie and Wager synthetic data experiments
The second synthetic data experiments use the setups of Nie and Wager [2021]. Four setups are used: A, B, C, and D. Each setup uses 5000 samples and is split into a training and testing set using a 50% split. Every setup is simulated with uncorrelated noise terms, uses five covariates, and a standard deviation of the error term of 1. Additionally, we make for each setup a heteroscedastic setting that makes the variance of the noise distribution dependent on the covariates, $\sigma=\sigma_{het}$. The same approach is taken here as in Section G.2.1 with 100 simulations, each using different random seeds and using the same $\alpha$ values. The CCT and CMC models also use normalized split WCPS. The results are shown in Figure 4, 5, 6, 7, and 16.
Besides the aforementioned experiments with the data-generating processes proposed by Nie and Wager [2021], we also perform a noise dependency impact experiment with it. This experiment will investigate the impact of changing the dependency between the noise variables, $\epsilon(0)$ and $\epsilon(1)$, of $Y(0)$ and $Y(1)$, respectively, on the different conformal frameworks. The only difference with the previous experiment is that the noise correlation coefficient $c$ changes from $-1$ to $1$ in increments of $0.1$. The results are shown in Figure 11, 15, and 14.
@@ G.2.2 Nie and Wager synthetic data experiments
Setup A simulates a difficult nuisance component and an easy treatment effect. $d$ is the number of covariates and $n$ the number of samples. $X$ is created as an independent variable, $n$ rows and $d$ columns thus consisting of $[X_{0},X_{1},...,X_{d-1}]$ each with length $n$. The other variables are the expected outcome $b$, the propensity of receiving treatment $\pi$, $\tau$ the CATE, $W$ the treatment assignment, $Y(0)$ outcome for control, $Y(1)$ for treatment, $\epsilon(0)$ noise variable for control, $\epsilon(1)$ noise variable for treatment, and $Y$ the true outcome, all with length $n$.
$X\sim Uniform([0,1],size=(n,d))$ (36)
$b=sin(\pi X_{0}X_{1})+2(X_{2}-0.5)^{2}+X_{3}+0.5X_{4}$ (37)
$\pi=max(0.1,min(sin(\pi X_{0}X_{1}),0.9)$ (38)
@@
$W\sim Binomial(1,\pi)$ (40)
$\sigma_{het}=\sigma*\log(1+|X_{1}|)$ (41)
$\epsilon(0)\sim\mathcal{N}(0,\sigma^{2})$ (42)
$\epsilon(1)=c\epsilon(0)+(1-c)\mathcal{N}(0,\sigma^{2})$ (43)
$Y(0)=b-0.5\tau+\epsilon(0)$ (44)
$Y(1)=b+0.5\tau+\epsilon(1)$ (45)
$Y=Y(0)(1-W)+Y(1)W$ (46)
@@
Setup D mimics a situation with unrelated treatment and control groups. All other variables are calculated similarly as in setup B except for $b$, $\pi$, and $\tau$.
@@
G.3 Semi-synthetic experiments
The learners are also evaluated on semi-synthetic datasets. Four semi-synthetic datasets/experiments are used: IHDP [Hill, 2011], NLSM [Carvalho et al., 2019], EDU [Zhou et al., 2021], and the ACIC 2016[Dorie et al., 2019]. The simulations run 100 times with $\alpha=0.1$, 100 MC samples, and with split WCPS using the residual as a conformity measure. Since these datasets do not provide propensity scores, we perform logistic regression to estimate them. For the ACIC 2016 experiments, we run the simulations 10 times because it has 77 different settings. The results are found in Figure 2, 8, 9, 12, and 17.
G.3.1 IHDP
The infant health and development program (IHDP) is a randomized controlled trial to evaluate the effect of home visits from physicians on the cognitive test scores of premature infants Hill [2011]. This dataset has 25 covariates with 6 continuous and 19 binary. Using this dataset, a semi-synthetic benchmark was created analogous to the study by Hill [2011], using setup B of their work. The used datasets are provided by Shalit et al. [2017]. The final dataset consists of 747 samples, 139 are treated samples and 608 of them are control samples. The data generation process of the semi-synthetic data is as follows:
$Y(0)\sim\mathcal{N}(exp(X+L)\beta,1)$ (57)
$Y(1)\sim\mathcal{N}(exp(X\beta)-\omega,1)$ (58)
@@ G.3.1 IHDP
respectively. $X$ represents the covariates in the IHDP dataset and $L$ is an offset matrix. The same treatment variable is kept as in the original IHDP study.
G.3.2 NLSM
The national study of learning mindsets (NLSM) is a large-scale randomized trial to investigate the effect of behavioural intervention on the academic performance of students [Carvalho et al., 2019]. Using the original NLSM dataset, new data is synthetically generated that follow the same distributions as the NLSM dataset as elaborated in Carvalho et al. [2019], Alaa et al. [2023]. The resulting dataset contains 10000 data points and 11 covariates ($X=[X_{0},X_{1},...,X_{10}]$). The data generation process for the semi-synthetic data is as follows:
@@ G.3.2 NLSM
$\tau=0.228+0.05\mathbb{I}(X_{6}<0.07)-0.05\mathbb{I}(X_{7}<-0.69)-0.08\mathbb{I}(X_{4}\in[1,13,14])+\gamma_{X_{0}-1}$ (60)
with $\mathbb{I}$ the indicator function and $\gamma_{X_{0}-1}$ the $(X_{0}-1)$th element of an array of Gaussian samples with size 76. Then the outcome is defined as follows:
$Y=W(Y-\tau)+(1-W)\tau$ (61)
@@
G.3.3 EDU
The EDU experimental setup, introduced by Zhou et al. [2022], evaluates the impact of maternal adult education on children’s learning outcomes using a semi-synthetic dataset derived from a randomized field experiment in India [Banerji et al., 2017]. The binary treatment corresponds to maternal education receipt, and the outcome is the difference in children’s test scores. After preprocessing (sample size: 8,627; 32 covariates), synthetic potential outcomes were generated by training neural networks $\hat{f}_{Y(0)}(\cdot)$ (control) and $\hat{f}_{Y(1)}(\cdot)$ (treatment) on observed outcomes. Uncertainty was modeled as $Y_{i}(0)\sim\hat{f}_{Y(0)}(X_{i})+(2-m_{i})\mathcal{N}(0,0.5^{2})$ and $Y_{i}(1)\sim\hat{f}_{Y(1)}(X_{i})+(2-m_{i})\exp(2)$, where $m_{i}$ indicates prior maternal education, introducing heterogeneous variance. Treatment imbalance was induced via a high-coefficient logistic propensity model $\mathbb{P}(W_{i}=1|X_{i})=(1+\exp(-X_{i}^{T}\beta))^{-1}$ and truncation of subjects with propensities in $[0.3,0.7]$. This design tests models under distributional heterogeneity and imbalance while aligning with real-world decision-making contexts.
G.3.4 ACIC 2016
The dataset from the 2016 Atlantic causal inference conference was created for a causal inference competition [Dorie et al., 2019]. It includes 4,802 observations with 58 covariates (3 categorical, 5 binary, 27 counts, and 23 continuous). The competition used these covariates to generate 77 different settings, each with varying outcome surface complexity, degrees of confounding, overlap, and treatment effect heterogeneity. We standardized all covariates and, for each setting, conducted 10 simulation runs, with a test set comprising 20% of the data, randomly generated, following the methodology of Curth and van der Schaar [2021], Alaa et al. [2023].